# Lightweight and Interpretable Transformer via Mixed Graph Algorithm Unrolling for Traffic Forecast

### v2

@@ 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in eq. 6 means
${\boldsymbol{\phi}}^{\tau+1}=$ $\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})=\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\bm{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}.$ (15)
We can compute eq. 15 element-wise as follows (see Appendix D for a derivation):
$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau}$ (16)
@@
Interpretation: Soft-thresholding in eq. 16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$. Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (Ortega et al., 2018), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli & Kovacevic, 2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$. Using eq. 16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved. For example, the constant (low-frequency) signal $\mathbf{1}$} passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$. Thus, attenuation of the high-pass channel means eq. 16 is a low-pass filter.
@@ 3.4.3 Updating Lagrange Multipliers
For objective eq. 6, we follow standard ADMM procedure to update Lagrange multiplier ${\bm{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
${\bm{\gamma}}^{\tau+1}$ $={\bm{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\bm{\gamma}}_{u}^{\tau+1}$ and ${\bm{\gamma}}_{d}^{\tau+1}$ in objective eq. 9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
${\bm{\gamma}}_{u}^{\tau+1}={\bm{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),~~~~~~~~~{\bm{\gamma}}_{d}^{\tau+1}={\bm{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).$ (18)
@@ 4 Unrolled Neural Network
We unroll our iterative ADMM algorithm to an interpretable neural net. The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (Bahdanau et al., 2014).
@@ 4.1 ADMM Algorithm Unrolling
We implement each iteration of our algorithm in Section 3.4 as a neural layer. Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems eq. 10 for ${\mathbf{x}}^{\tau+1}$, eq. 11 for ${\mathbf{z}}_{u}^{\tau+1}$, and eq. 12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via eq. 16. CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E for detailed implementation). Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer. Multipliers ${\bm{\gamma}}^{\tau+1},{\bm{\gamma}}_{u}^{\tau+1},{\bm{\gamma}}_{d}^{\tau+1}$ are then updated via eq. 17 and eq. 18 to complete one layer. Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\bm{\Theta}}_{b}$. See Fig. 3 for an illustration.
@@ 4.2 Self-Attention Operator in Transformer
We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (Bahdanau et al., 2014). Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$. The affinity between tokens $i$ and $j$, $e(i,j)$, is the dot product of linear-transformed ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{Q}},{\mathbf{K}}\in\mathbb{R}^{E\times E}$ are the query and key matrices, respectively. Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
$a_{i,j}=\frac{\exp(e(i,j))}{\sum_{l=1}^{N}\exp(e(i,l))},~~~~~~e(i,j)=({\mathbf{Q}}{\mathbf{x}}_{j})^{\top}({\mathbf{K}}{\mathbf{x}}_{i}).$ (19)
@@
${\mathbf{y}}_{i}=\sum_{l=1}^{N}a_{i,l}{\mathbf{x}}_{l}{\mathbf{V}}$ (20)
@@ 4.3 Graph Learning Modules
Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$. For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in Feng & Tassiulas (2022), and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details). As done in Thuc et al. (2024), the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
$\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})$ (21)
where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$. Distance in eq. 21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$. Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
$w^{u}_{i,j}=\frac{\exp\left(-\text{d}^{u}(i,j)\right)}{\sqrt{\sum_{l\in{\mathcal{N}}_{i}}\exp(-\text{d}^{u}(i,l))}\sqrt{\sum_{k\in{\mathcal{N}}_{j}}\exp(-\text{d}^{u}(k,j))}}$ (22)
@@
Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in eq. 22 are essentially attention weights $a_{i,j}$’s in eq. 19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}. Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—eq. 10, eq. 11, eq. 12 and eq. 16—derived from our optimization of objective eq. 4.
Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$. For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$. Like eq. 21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}. Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization, as done in eq. 22 for an undirected edge.
Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\bm{\Phi}}_{b}$.
@@ 5.1 Experimental Setup
We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (Li et al., 2017) and traffic flow dataset PEMS03 (Guo et al., 2022), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors. We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly, and split it into training, validation, and test sets by 6:2:2. We predict the traffic speed/flow in the following 30 / 60 /120 minutes (6 / 12 /24 steps) from the previously observed 60 minutes (12 steps) for all datasets. We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence. For undirected graphs ${\mathcal{G}}^{u}$, we connect each node to its $k$ nearest neighbors (NNs). We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a pair of graph learning modules before each ADMM block. We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism. We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths. See Appendix F for more details.
@@ 5.2 Experimental Results
Table 1: Comparison of categorized baselines and model parameters for 60-minute forecasting on the PEMS03 dataset against ours.
Category Selected Models Params #
Model-based VAR (Reinsel, 2003) -
[…]
GATs GMAN (Zheng et al., 2020), ST-Wave (Fang et al., 2023) 210K, 883K
Transformers PDFormer (Jiang et al., 2023), STAEformer (Liu et al., 2023) 531K, 1404K
Adaptive-graph models Graph WaveNet (GWN) (Wu et al., 2019), AGCRN (Bai et al., 2020) 277K, 749K
Simple linear models STID (Shao et al., 2022), SimpleTM (Chen et al., 2025) 123K, 540K
Mixed-graph unrolling Ours 34K
  • $\dagger$
    STGCN predicts only one time step at a time. We train the model to predict only the next step, and predict the entire sequence recursively.
@@
Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1315 edges)
@@
Ours 26.10 / 17.03 / 18.85 27.67 / 17.46 / 17.72 34.40 / 21.16 / 23.47 10.35 / 4.22 / 9.70 12.34 / 5.18 / 11.80 15.19 / 7.56 / 16.12
@@
Table 2 shows the performance of our models and baselines in traffic forecasting, and the last column of Table 1 compares the number of parameters. We observe that our models reach comparable performance to most baseline models, while employ drastically fewer parameters (6.4% of transformer-based PDFormer). Among all the models, ST-Wave and SimpleTM generally achieve the best performance. However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (Battaglia et al., 2018); both models require even more parameters than PDFormer. Our unrolled model also generally outperforms the multi-attention model GMAN. Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance. The largest GCN model, STSGCN, fails to achieve a satisfying performance for PEMS03, where the real connections are highly sparse. STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 datasets. In summary, none of the DL-based baseline schemes unambiguously achieves the best performance across the two datasets, while all require significantly more parameters than our models. In contrast, our model captures the complex and evolving spatial-temporal relationship accurately using a mixed graph, demonstrating consistent performance for all predicting horizons; notably, the performance gap to the best model decreases as the horizon grows.
Further experiments on data efficiency are presented in Appendix G, and a comprehensive ablation study is provided in Appendix H.

### v3

@@ 3.4.2 Minimizing ${\boldsymbol{\phi}}^{\tau+1}$
Fixing ${\mathbf{x}}^{\tau+1}$ at iteration $\tau$, optimizing ${\boldsymbol{\phi}}^{\tau+1}$ in Equation 6 means
$~\textstyle{{\boldsymbol{\phi}}^{\tau+1}=\arg\min_{{\boldsymbol{\phi}}}g({\boldsymbol{\phi}})}$ (15)
$=$ $~\textstyle{\mu_{d,1}\|{\boldsymbol{\phi}}\|_{1}+({\boldsymbol{\gamma}}^{\tau})^{\top}({\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1})+\frac{\rho}{2}\|{\boldsymbol{\phi}}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}\|^{2}_{2}}.$
We can compute Equation 15 element-wise as follows (see Appendix D for a derivation):
$\delta$ $=({\mathbf{L}}^{d}_{r})_{i}{\mathbf{x}}^{\tau+1}-\rho^{-1}\gamma_{i}^{\tau},$ (16)
@@
Interpretation: Soft-thresholding in Equation 16 to compute $\phi_{i}^{\tau+1}$ attenuates the $i$-th entry of ${\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}$. Given that graph Laplacians—second difference matrices on graphs—are high-pass filters (ortega18ieee), we interpret ${\mathbf{L}}^{d}_{r}$ as the high-pass channel of a two-channel filterbank (Vetterli2013), where the corresponding low-pass channel is ${\mathbf{I}}-{\mathbf{L}}^{d}_{r}={\mathbf{W}}^{d}_{r}$. Using Equation 16 for processing means that the low-pass channel ${\mathbf{W}}^{d}_{r}$ is unused and thus preserved. For example, the constant (low-frequency) signal $\mathbf{1}$} passes through the low-pass channel undisturbed, i.e., ${\mathbf{1}}={\mathbf{W}}^{d}_{r}{\mathbf{1}}$, while it is filtered out completely by the high-pass channel, i.e., ${\mathbf{0}}={\mathbf{L}}^{d}_{r}{\mathbf{1}}$. Thus, attenuation of the high-pass channel means Equation 16 is a low-pass filter.
@@ 3.4.3 Updating Lagrange Multipliers
For objective Equation 6, we follow standard ADMM procedure to update Lagrange multiplier ${\boldsymbol{\gamma}}^{\tau+1}$ after each computation of ${\mathbf{x}}^{\tau+1}$ and ${\boldsymbol{\phi}}^{\tau+1}$:
${\boldsymbol{\gamma}}^{\tau+1}$ $={\boldsymbol{\gamma}}^{\tau}+\rho({\boldsymbol{\phi}}^{\tau+1}-{\mathbf{L}}^{d}_{r}{\mathbf{x}}^{\tau+1}).$ (17)
Splitting the $\ell_{2}$-norm terms means that additional multipliers ${\boldsymbol{\gamma}}_{u}^{\tau+1}$ and ${\boldsymbol{\gamma}}_{d}^{\tau+1}$ in objective Equation 9 must also be updated after each computation of ${\mathbf{x}}^{\tau+1}$, ${\mathbf{z}}_{u}^{\tau+1}$, ${\mathbf{z}}_{d}^{\tau+1}$, and ${\boldsymbol{\phi}}^{\tau+1}$:
$\begin{aligned} {\boldsymbol{\gamma}}_{u}^{\tau+1}&={\boldsymbol{\gamma}}_{u}^{\tau}+\rho_{u}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{u}^{\tau+1}),\\ {\boldsymbol{\gamma}}_{d}^{\tau+1}&={\boldsymbol{\gamma}}_{d}^{\tau}+\rho_{d}({\mathbf{x}}^{\tau+1}-{\mathbf{z}}_{d}^{\tau+1}).\end{aligned}.$ (18)
@@ 4 Unrolled Neural Network
We unroll our iterative ADMM algorithm to an interpretable neural net. The crux to our network is two periodically inserted graph learning modules for undirected and directed edge weight computation that are akin to the classical self-attention mechanism in transformers (bahdanau14).
@@ 4.1 ADMM Algorithm Unrolling
We implement each iteration of our algorithm in Section 3.4 as a neural layer. Specifically, each ADMM iteration is implemented as four sub-layers in sequence: solving linear systems Equation 10 for ${\mathbf{x}}^{\tau+1}$, Equation 11 for ${\mathbf{z}}_{u}^{\tau+1}$, and Equation 12 for ${\mathbf{z}}_{d}^{\tau+1}$ via CG, and computing ${\boldsymbol{\phi}}^{\tau+1}$ via Equation 16. CG is itself an iterative descent algorithm, where parameters $\alpha$ and $\beta$ corresponding to step size and momentum can be tuned end-to-end (see Appendix E.1 for detailed implementation). Weight parameters $\mu_{u},\mu_{d,2},\mu_{d,1}$ for prior terms and ADMM parameters $\rho,\rho_{u},\rho_{d}$ are also tuned per layer. Multipliers ${\boldsymbol{\gamma}}^{\tau+1},{\boldsymbol{\gamma}}_{u}^{\tau+1},{\boldsymbol{\gamma}}_{d}^{\tau+1}$ are then updated via Equation 17 and Equation 18 to complete one layer. Parameters learned during back-propagation for each ADMM block $b$ are denoted by ${\boldsymbol{\Theta}}_{b}$. See Fig. 3 for an illustration, and Appendix E.2 for the detailed algorithms.
@@ 4.2 Self-Attention Operator in Transformer
We first review the basic self-attention mechanism: a scaled dot product following by a softmax operation (bahdanau14). Denote by ${\mathbf{x}}_{i}\in\mathbb{R}^{E}$ an embedding for token $i$. The affinity between tokens $i$ and $j$, $e(i,j)$, is the dot product of linear-transformed ${\mathbf{K}}{\mathbf{x}}_{i}$ and ${\mathbf{Q}}{\mathbf{x}}_{j}$, where ${\mathbf{Q}},{\mathbf{K}}\in\mathbb{R}^{E\times E}$ are the query and key matrices, respectively. Using softmax, non-negative attention weight $a_{i,j}$ is computed from $e(i,j)$’s as
$\textstyle{a_{i,j}=\frac{\exp(e(i,j))}{\sum_{l=1}^{N}\exp(e(i,l))},~~e(i,j)=({\mathbf{Q}}{\mathbf{x}}_{j})^{\top}({\mathbf{K}}{\mathbf{x}}_{i})}.$ (19)
@@
$\textstyle{{\mathbf{y}}_{i}=\sum_{l=1}^{N}a_{i,l}{\mathbf{x}}_{l}{\mathbf{V}}},$ (20)
@@ 4.3 Graph Learning Modules
Undirected Graph Learning: We learn an undirected graph ${\mathcal{G}}^{u}$ in a module $\texttt{UGL}_{b}$ at block $b$. For each node $i$, we compute a low-dimensional feature vector ${\mathbf{f}}_{i}^{u}=F^{u}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$, where ${\mathbf{e}}_{i}\in\mathbb{R}^{E}$ is a vector of signal values, position embeddings in time and Laplacian embeddings of the physical graph as done in feng2022, and $F^{u}(\cdot)$ is a chosen learned nonlinear feature function $F^{u}:\mathbb{R}^{E}\mapsto\mathbb{R}^{K}$, e.g., a shallow GNN (see Appendix F.3 for details). As done in Do2024, the feature distance between nodes $i$ and $j$ is computed as the Mahalanobis distance $\text{d}^{u}(i,j)$:
$\textstyle{\text{d}^{u}(i,j)=({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})^{\top}{\mathbf{M}}({\mathbf{f}}_{i}^{u}-{\mathbf{f}}_{j}^{u})},$ (21)
where ${\mathbf{M}}\succeq 0$ is a symmetric PSD metric matrix of dimension $\mathbb{R}^{K\times K}$. Distance in Equation 21 is non-negative and symmetric, i.e., $\text{d}^{u}(i,j)=\text{d}^{u}(j,i)$. Weight $w^{u}_{i,j}$ of undirected edge $(i,j)\in{\mathcal{E}}^{u}$ is computed as
$\textstyle{w^{u}_{i,j}=\frac{\exp\left(-\text{d}^{u}(i,j)\right)}{\sqrt{\sum_{l\in{\mathcal{N}}_{i}}\exp(-\text{d}^{u}(i,l))}\sqrt{\sum_{k\in{\mathcal{N}}_{j}}\exp(-\text{d}^{u}(k,j))}}},$ (22)
@@
Remark: Interpreting distance $\text{d}^{u}(i,j)$ as e(i,j)-e(i,j), edge weights $w^{u}_{i,j}$’s in Equation 22 are essentially attention weights $a_{i,j}$’s in Equation 19, and thus the undirected graph learning module constitutes a self-attention mechanism, albeit requiring fewer parameters, since parameters for function $F^{u}(\cdot)$ and metric matrix $\mathbf{M}$} can be much smaller than large and dense query and key matrices, $\mathbf{Q}$} and $\mathbf{K}$}. Further, no value matrix $\mathbf{V}$} is required, since output signal ${\mathbf{x}}^{\tau+1}$ is computed through a set of low-pass filters—Equation 10, Equation 11, Equation 12 and Equation 16—derived from our optimization of objective Equation 4.
Directed Graph Learning: Similarly, we learn a directed graph ${\mathcal{G}}^{d}$ in a module $\texttt{DGL}_{b}$ at block $b$. For each node $i$, we compute a feature vector ${\mathbf{f}}^{d}_{i}=F^{d}({\mathbf{e}}_{i})\in\mathbb{R}^{K}$ using feature function $F^{d}(\cdot)$. Like Equation 21, the Mahalanobis distance $\text{d}^{d}(i,j)$ between nodes $i$ and $j$ is computed from ${\mathbf{f}}_{i}^{d}$ and ${\mathbf{f}}_{j}^{d}$ with a learned PSD metric $\mathbf{P}$}. Weight $w_{j,i}^{d}$ of directed edge $i,j]\in{\mathcal{E}}^{d}$ is then computed using exponentials and normalization as
${w_{j,i}^{d}=\frac{\exp(-\text{d}^{d}(j,i)}{\sum_{(i,k)\in{\mathcal{E}}^{d}}\exp(-\text{d}^{d}(k,i))}.}$ (23)
Parameters learned via back-propagation by the two graph learning modules at block bb—parameters of $F^{u}(\cdot)$, $F^{d}(\cdot)$, $\mathbf{M}$} and $\mathbf{P}$}—are denoted by ${\boldsymbol{\Phi}}_{b}$.
@@ 5.1 Experimental Setup
We evaluate our model’s performance on commonly used traffic speed dataset METR-LA (li2017diffusion) and traffic flow dataset PEMS03 (9346058), each of which contains city-size traffic data with a sampling interval of 5 minutes in a local region, together with the real connections and distances (or travel cost) between sensors. We reduce the size of the original dataset by sub-sampling 1/3 of all data uniformly [^8] With a drastically smaller parameter count, our model focuses on the efficiency in utilizing the observations and the effectiveness with small training sets. Uniform subsampling reduces the size of the training set while making full use of the entire observation. See a detailed analysis in Appendix G.4., and split it into training, validation, and test sets by 6:2:2. We predict the traffic speed/flow in the following 30 / 60 /120 minutes (6 / 12 /24 steps) from the previously observed 60 minutes (12 steps) for all datasets. We restrict our model to learn a sparse graph based on the real road connections and a limited time window to model only the local spatial/temporal influence. For undirected graphs ${\mathcal{G}}^{u}$, we connect each node to its $k$ nearest neighbors (NNs). We construct 5 ADMM blocks, each contains 25 ADMM layers, and insert a pair of graph learning modules before each ADMM block. We set feature dimensions $K=6$, and stack vertically 4 graph learning modules to learn in parallel as an implementation of the multi-head attention mechanism. We minimize the Huber loss with $\delta=1$ for the entire reconstructed sequence and the groundtruths. See Appendix F for more details.
@@ 5.2 Experimental Results
Dataset & PEMS03 (358 nodes, 547 edges) METR-LA (207 nodes, 1,315 edges)
@@
Ours 25.05 / 15.85 / 16.49 26.96 / 16.58 / 18.07 34.06 / 20.23 / 23.86 10.06 / 4.05 / 9.23 12.17 / 5.27 / 11.78 15.19 / 7.14 / 16.44
Table 2: Comparison of categorized baselines and model size for 60-minute forecasting on the PEMS03 dataset against ours.
Category Selected Models Params #
Model-based VAR (reinsel2003elements) -
[…]
STSGCN (song2020spatial) 3,496K
GATs GMAN (zheng2020gman) 210K
ST-Wave (fang2023spatio) 883K
Transformers PDFormer (jiang2023pdformer) 531K
STAEformer (liu2023spatio) 1,404K
Adaptive-graph models Graph WaveNet (10.5555/3367243.3367303) 277K
AGCRN (10.5555/3495724.3497218) 749K
MLP-based models STID (10.1145/3511808.3557702) 123K
SimpleTM (chen2025simpletm) 540K
Mixed-graph unrolling Ours 38K
@@
Table 1 shows the performance of our models and baselines in traffic forecasting, and the last column of Table 2 compares the number of parameters. We observe that our models achieve comparable performance to most baseline models, achieving top-3 performance in at least one metric for 30- and 60-minute forecasting, while employing drastically fewer parameters (7.2% of transformer-based PDFormer). Among all the models, ST-Wave and SimpleTM generally achieve the best performance. However, ST-Wave relies on an application-specific usage of wavelet transforms, and SimpleTM employs linear models with no inductive bias, which is parameter-heavy (47094) and lacks interpretability; both models require even more parameters than PDFormer. The largest GCN model, STSGCN, fails to achieve a satisfying performance for PEMS03, where the real connections are highly sparse. STAEformer, the largest transformer, also overfits heavily in the reduced PEMS03 datasets. We also observe that our unrolled model also generally outperforms the multi-attention model GMAN. Though both our model and GMAN leverage local graph neighborhoods when computing attention, we dramatically reduce parameters by replacing classical self-attention mechanisms in transformers with unrolled ADMM iterations with no observable degradation in performance. In summary, no single DL-based baseline consistently outperforms the others across both datasets; furthermore, all baselines require significantly more parameters than our model. In contrast, our model achieves top-3 forecasting performance in two of three prediction windows for both datasets (top-5 in the remaining prediction window), in at least one metric.
We plot the three metrics against parameter size (with computational cost coded in color) for 60-minute forecast on the PEMS03 datasets in Figure 4. The conventional baseline models with larger parameter counts generally exhibit lower prediction errors, showing an overall downward trend in the plot, except the outlier STAEFormer which overfits heavily. In contrast, our model lies distinctly apart from the main trend in the lower-left corner in each plot, delivering the accuracy of models over $10\times$ larger while using dramatically fewer parameters.
Figure 4: The trade-off between the number of parameters and performance for 60-minute forecast on PEMS03 dataset. The color of each scatter point represents the inference computational cost (in MFLOPs) in Table 7 in Appendix G.3.
Additional experiments are presented in Appendix G.
5.3 Ablation Studies
For ablation studies, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, and the benefit of directed-graph modeling of temporal relations.
First, we discuss the effectiveness of the new directed graph priors, DGLR and DGTV, as well as the benefit of directed-graph modeling of temporal relations. We separately test the effects of DGLR and DGTV by removing the prior and unrolling the ADMM algorithm. We denote these models as “w/o DGTV” and “w/o DGLR”. To test the benefit of the directed-graph modeling, we change the temporal graph ${\mathcal{G}}^{d}$ into a fully undirected graph ${\mathcal{G}}^{n}$ while maintaining the connections, and replace the directed graph priors with a GLR prior ${\mathbf{x}}^{\top}{\mathbf{L}}^{n}{\mathbf{x}}$ with the same coefficient $\mu_{n}=\mu_{d,2}$. Detailed derivations of the ADMM algorithm for all the above ablation studies are in Appendix I.
Additionally, we ran ablations of the splitting-term strategy by optimizing ${\mathbf{x}}^{\tau+1}$ using only Equation 7. However, the unrolled model crashed within the first few training epochs due to numerical instability, indicating that splitting the $\ell_{2}$ terms can stabilize the training process, since CG in solving Equation 7 converges much slower than in solving Equation 10-Equation 12. See Appendix H for further ablation studies in the number of nearest neighbors $k$, time window size $W$,and number of ADMM blocks and layers.
Table 3: Performance comparison in RMSE / MAE / MAPE(%) of the ablation studies for mixed-graph modeling and directed graph priors in 60-minute forecasting with PEMS03 and METR-LA.
Model PEMS03 METR-LA
Undirected graph 29.00 / 18.21 / 19.99 12.53 / 5.37 / 11.97
w/o DGTV 27.62 / 17.54 / 19.04 12.59 / 5.39 / 12.26
w/o DGLR 30.53 / 18.69 / 19.06 12.45 / 5.34 / 11.98
Ours (Mixed graph) 26.96 / 16.58 / 18.07 12.17 / 5.27 / 11.78
5.4 Interpreting the Learned Graph via Eigenvector Centrality
We analyze the relationship between the last undirected graph learned by the model and the observed traffic flows. For the best trained model on a 60-minute forecast on PEMS03, we calculate the learned weighted adjacency matrices ${\mathbf{W}}^{u}$ of the undirected graphs on the first 288 time steps (24 hours) in the test set. We calculate the normalized positive eigenvector corresponding to the largest eigenvalue of ${\mathbf{W}}^{u}$, known as the Perron vector, which is strictly positive by the Perron-Frobenius Theorem (johnson12). The Perron vector ${\mathbf{v}}\in\mathbb{R}^{N}$ reflects the eigenvector centrality (bonacich87) of the individual nodes: the “importance” of individual nodes, where a node is deemed important if it is connected to other important nodes, i.e., $\lambda v_{i}=\sum_{j}W^{u}_{i,j}v_{j}$ for $v_{j}>0$.
(a) Neighborhood of nodes #27 and #350.
(b) Neighborhood of node #171.
(c) Neighborhood of node #328.
Figure 5: The traffic patterns of the nodes in the neighborhood of the selected important nodes by Perron vectors. The traffic flows of selected nodes are bolded.
Figure 6: The evolution of eigenvector centrality in the learned undirected graph in the last block and traffic flow in the first 24 hours in the test split of the PEMS03 dataset.
Figure 6 shows the evolution of the top-10 Perron vector entries and the corresponding traffic flow. First, we observe that evident disturbances in the Perron vector entries only exist when there is a rapid lift in the traffic flows from near zero to hundreds, practically corresponding to the beginning of a new morning. Furthermore, the eigenvector centrality returns to the initial state once the traffic flow becomes stable, and there are no significant changes in the ranking of the eigenvector centrality, indicating that the node set with the largest Perron vector entries is generally consistent throughout the observation, which marks the critical congestion hubs.
To further examine the role of the discovered important nodes in short-range interactions, we then select the four most important nodes #328, #27, #350 and #171 according to the elements in the Perron vectors, and plot the traffic flow of each node and the physical neighbors in selected windows of 30 and 60 steps in the daytime in Figure 5, where the selected important nodes are plotted with bolder lines (note that the neighborhood of nodes #27 and #350 are exactly the same set). In the red box in Figure 5(a), the spike in node #27 results in the negative spike in the neighboring nodes #274, #24 and #351. Further, the combined effect of this spike and the bump in node #350 results in the overshoot after the downward spike in node #274. The signal trends in the red box in Figure 5(b) clearly reflect the time delay of the traffic flow in the neighboring nodes except the “quiet sequences” to the selected node #171. Both figures 5(a) and 5(b) demonstrate clear spatiotemporal dependency patterns captured by the learned undirected graph, where high-centrality nodes play a dominant role in propagating local disturbances and temporal delays to their physical neighbors. Additionally, as shown in Figure 5(c), for the traffic patterns in the neighborhood of node #328, which has the largest Perron vector entry, two of its three neighboring nodes exhibit the 5th (#171) and 7th (#167) largest eigenvector centralities. This observation is consistent with the rationale that the most important node tends to be connected to other highly important nodes. Consequently, the traffic flow trends in the neighborhood of node #350 tend to be less affected by the central node, as its neighbors exhibit relatively high eigenvector centralities themselves.
In conclusion, our model successfully discovers physically interpretable graph structures from data, and the learned node importance is highly consistent with actual traffic dynamics. Practically, our model can identify the key intersections that critically influence traffic dynamics and further guide traffic engineering in alleviating future traffic congestion.